OpenStudy (anonymous):

A 10−pound package of cookies costs (y + 1) dollars while a 6−pound package costs (y − 1) dollars. If the price per pound of cookies is the same for both packages, the equation below can be used to solve for y.

OpenStudy (anonymous):

\[\frac{ y+1 }{ 10 } = \frac{ y-1 }{ 6 }\]

OpenStudy (anonymous):

6(y+1)=10(y-1) 6y+6 = 10y-10 4y=16 y=?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Divide both sides by 4.

OpenStudy (texaschic101):

4y = 16 4y/4 = 16/4 y = ?

OpenStudy (anonymous):

@Albany_Goon, what is 16 divided by 4?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Yes, so y=4

OpenStudy (anonymous):

thnx are u up for more

OpenStudy (anonymous):

I'll try.

OpenStudy (anonymous):

A greenhouse that specializes in growing geraniums is divided into sections. The number of geranium pots in each section depends on the number of sprinklers in that section. There are x(7x − 2) pots of red geraniums and (x + 3)^2 pots of pink geraniums in a section with x sprinklers. Which simplified expression represents the total number of red and pink geranium pots in a section with x sprinklers?

OpenStudy (anonymous):

@texaschic101, do you want to take this one?

OpenStudy (anonymous):

HELPPP

OpenStudy (anonymous):

@factor or @CO_oLBoY

OpenStudy (anonymous):

@hba @wolfe8 @eliassaab

OpenStudy (anonymous):

i think it should be like this but u should confirm it x(7x − 2) + (x + 3)^2 7x^2 - 2x + x^2 + 9 + 6x 8x^2 + 4x + 9 i think this should be ur required expression. u can also factorize it

OpenStudy (anonymous):

yes thnx you

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

one more please

OpenStudy (anonymous):

ok but ask fast i dont have much time

OpenStudy (anonymous):

A news agent conducted a survey among travel magazine subscribers of a town and found that there was circulation of only three travel magazines. He made the Venn diagram below to show the number of subscribers to each of the three magazines: Leisure, Outbound, and Vacations. How many subscribers belong to the set V U (~O ∩ L)?

OpenStudy (anonymous):

choices 749 821 870 954

OpenStudy (anonymous):

sorry this set is like that V U (O ∩ L) = {84,49,49,72,67,387} so, 84+49+49+72+67+387=708 let me check there must some other mistake too

OpenStudy (anonymous):

can u check ur choices 1 more time

OpenStudy (anonymous):

those are the choices

OpenStudy (anonymous):

is this is asked V U (~O ∩ L)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then the ans should be 708

OpenStudy (anonymous):

ok thnx

OpenStudy (anonymous):

thats not one of them

OpenStudy (anonymous):

i know but any other doesnt make sense

OpenStudy (anonymous):

thnx anyway

OpenStudy (anonymous):

u r welcome